AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME  ELOG logo
Message ID: 291     Entry time: Sun Feb 27 00:53:24 2022
Author: David, Pierre-Michell 
Subject: Current status 

We have recovered the beam at 01:49, after such a long time we had no target so we increased from 80K to 120K at 01:52. 02:01 update 120K-->95K as we have observed around that temperature always fluctuations and is safer.

Ring current 7.74e7 particles/s

Merger, Tape server (WS3): 1619 kB/s writing ok

Leakage current (WS4): 3.26 uA

DAQ (WS2): temperature checked, clock status: ok, adc calibration: ok, white rabbit check: failed

Online monitoring (WS 5): done

 

Attachment 1: 2022-02-27_01-58-53xytotal.png  181 kB  | Hide | Hide all
2022-02-27_01-58-53xytotal.png
Attachment 2: 2022-02-27_01-58-34exrate.png  156 kB  | Hide | Hide all
2022-02-27_01-58-34exrate.png
Attachment 3: 2022-02-27_01-58-18extotal.png  165 kB  | Hide | Hide all
2022-02-27_01-58-18extotal.png
Attachment 4: 2022-02-27_01-57-56current.png  20 kB  | Hide | Hide all
2022-02-27_01-57-56current.png
Attachment 5: 2022-02-27_01-57-38goodevents.png  103 kB  | Hide | Hide all
2022-02-27_01-57-38goodevents.png
Attachment 6: 2022-02-27_01-57-16temp.png  111 kB  | Hide | Hide all
2022-02-27_01-57-16temp.png
Attachment 7: 2022-02-27_01-57-03aida.png  165 kB  | Hide | Hide all
2022-02-27_01-57-03aida.png
ELOG V3.1.4-unknown